Archive of Monthly reports
Day (DOY) Temp (C) Width (DAY) 294.7 -116.1 0.28 297.3 -116.5 0.33 299.9 -117.6 0.28 Small, brief spike to -119.0 at Day 300.9 due to ACIS warm reboot after DPA anomaly (see Notes)
10/23-24 TSCIUSF1 S/C Temp SC-IUS FITTING-1 TEMP reaches upper yellow limit 349.7
10/26 DPA A off - See Notes below - Lost obsids 979 980
|MSID||DESCRIPTION||COMMENTS||5EIOT||EIO TEMP||- climbed steadily up 15 degrees K since launch||5EPHINT||EPHIN TEMP||- climbed steadily up 15 degrees K since launch||HKABIASLEAKI||HK A bias leakage current||- climbed steadily up 0.2 amp since launch||HKBBIASLEAKI||HK B bias leakage current||- climbed steadily up 0.1 amp since launch||HKCBIASLEAKI||HK C bias leakage current||- climbed steadily up 2 amp since launch||HKDBIASLEAKI||HK D bias leakage current||- climbed steadily up 2 amp since launch||HKEBIASLEAKI||HK E bias leakage current||- climbed steadily up 2 amp since launch||HKFBIASLEAKI||HK F bias leakage current||- rate of rise seems to be increasing since August total clime 0.1 amp||HKEBOXTEMP||HK EBox tempreture||- climbed steadily up 12 degrees K since launch||TEIO||EPHIN ELECTRONICS HOUSING TEMP||- rate of rise seems to be decreasing up 15 degrees K since launch||TEPHIN||EPHIN SENSOR HOUSING TEMP||- rate of rise seems to be decreasing up 12 degrees K since launch||HKGHV||HK G high voltage||- dropped 1 ADU around January, flat otherwise||HKN6I||HK -6V rail current||- flat||HKN6V||HK -6V rail voltage||- rate of rise seems to be increasing total rise 0.4 V||HKP27I||HK +27V rail current||- climbed steadily up 0.02 V since launch||HKP27V||HK +27V rail voltage||- flat||HKP5I||HK +5V rail current||- flat||HKP5V||HK +5V rail voltage||- flat||HKP6I||HK +6V rail current||- flat||HKP6V||HK +6V rail voltage||- flat|
|MSID||DESCRIPTION||COMMENTS||SCCT0||Sci single detector counter CT0||- Low rate down from 6 to 2, perhaps rising again since day 400||SCE1300||Sci single detector counter E1300||- flat||SCE150||Sci single detector counter E150||- flat||SCE300||Sci single detector counter E300||- flat||SCE3000||Sci single detector counter E3000||- flat||SCH25GM||Sci single detector counter H25GM||- flat||SCH41GM||Sci single detector counter H41GM||- flat||SCH4GM||Sci single detector counter H4GM||- flat||SCH8GM||Sci single detector counter H8GM||- flat||SCINT||Sci single detector counter INT||-High rate down from 120 to 60, perhaps rising again since day 400||SCP25GM||Sci single detector counter P25GM||- flat||SCP41GM||Sci single detector counter P41GM||- flat||SCP4GM||Sci single detector counter P4GM||- flat||SCP8GM||Sci single detector counter P8GM||- flat|
|OBSID||DETECTOR||GRATING||TARGET||ANALYSIS||ACA||1032||ACIS-I||NONE||RX J0852.0-4622||OK||OK||434||ACIS-S||NONE||3C 212||Aspect prob?||OK||1636||ACIS-S||NONE||NGC 2683||OK||OK||1596||ACIS-S||NONE||0902+343||OK||OK||1635||ACIS-S||NONE||NGC 2541||OK||Missing||976||HRC-S||LETG||KPD0005+5106||OK||OK||1591||HRC-S||LETG||NGC 3227||OK||OK||975||HRC-S||LETG||AD LEO||OK||OK||755||ACIS||NONE||B2224+65||OK||OK||2161||ACIS-I||NONE||2MASSJ102724+1219||OK||OK||2200||ACIS||NONE||R1361||OK||PROBLEM||1659||ACIS-I||NONE||ZWCL 1953||OK||OK||552||ACIS-I||NONE||RX J2129.6+0005||OK||OK||314||ACIS||NONE||M31 TRANSIENT||OK||OK||894||ACIS-I||NONE||ABELL 2255||OK||OK||2146||ACIS-S||NONE||2251+134||OK||OK||1031||ACIS-S||NONE||Mrk 231||OK||OK||2199||ACIS-S||NONE||RXJ0819.6+6336||OK||OK||2202||ACIS-I||NONE||RBS797||OK||PROBLEM||1872||ACIS-S||NONE||NGC 2068||OK||OK|
From: Roger Brissenden
During last week the observing schedule was completed as planned with the exception of approximately 6 hours of ACIS observing that was impacted when the ACIS Digital Processing Assembly (DPA-A) powered off at 11:42 am EDT on 26 Sept. The turn-off was detected during a real-time pass and following engineering assessment, the DPA was powered on and ACIS returned to science operation. ACIS has operated nominally since then and the cause of the turn-off is being investigated by the ACIS engineering team. Observations of HD100546, HD104237 and a portion of Sgr A were impacted and will be rescheduled.
From: Paul Plucinsky
For DPA A we need to watch: 1DP28AVO - input voltage which is currently 29.92 V,normal 1DPICACU - DPA Input Current which is currently ZERO, definitely not normal 1DPP0AVO - DPA +5V Analog which is currently 0.02 V which is consistent with being off, not normal 1DPPSAX - DPA A Enabled/Disabled, currently enabled, normal 1DPPSA - DPA A On/Off, currently OFF !!! DPA B looks OK to me: 1DP28BVO - input voltage which is currently 30.08 V,normal 1DPIBCU - the input current on side B of the DPA currently 0.27 A, which is consistent with the FEPs on side B begin off. We always keep the FEPs on in the load. So, I am not sure why the FEPs on side B are off. Perhaps Bob can explain why cutting the power on side DPA A would cause the B side FEPs to turn off. But 0.3 A and 8.0 W is a reasonable value for he side B power consumption. 1DPP0BVO - DPA +5V Analog Side B which is currently 5.13 V which is consistent with DPA B being on. 1DPPSBX - DPA B Enabled/Disabled, currently enabled, normal 1DPPSB - DPA B On/Off, currently ON, normal SO, DPA A is enabled but off, DPA B looks fine. The DPA A Over-current indicator 1DPCPAOC indicates that we did not have an overcurrent situation.